# High-Precision Physics Experiments at Huizhou Large-Scale Scientific Facilities

### v1

@@ II Huizhou Hadron Spectrometer
HIAF will be able to provide intense proton beams up to 9.3 GeV and heavy-ion (e.g., uranium) beams up to 2.45 GeV/u. The CNUF is designed to accelerate uranium beams up to an energy of 9.1 GeV/u. This provides good opportunities for a wide range of physical research, including the search for new particles and new interactions beyond the standard model, the test of basic symmetries, the search for new (exotic) hadron states, exotic di-baryons and new (multi-strange) hypernuclei, precise measurement of hadron and hypernucleus prop- erties, the search and location of the nuclear matter phase boundary and critical point, etc. In order to conduct these physics studies, we plan to build an experimental spectrometer at the HIAF high energy terminal, called the Huizhou Hadron Spetrometer (HHaS). Successful construction of HHaS will promote the development of medium- and high-energy particle physics and nuclear physics researches based on facilities in China.
@@ II.1 Conceptual Design
The 5-dimensional tracking system features an all-silicon pixel detector with $\sim$100 μm pitch pixels, incorporating both energy deposition and time measurement capabilities on each pixel. With $\sim$<∼ <20-μm hit position resolution and a 1.6-Tesla magnetic field, $\sim$3momentum resolution can be achieved with a tracker radius of merely 30 cm. This makes HHaS a very compact and cost-effective spectrometer. The good hit resolution also provides exceptional vertex resolution, which is crucial for clear reconstruction of strange particles and hypernuclei. The 10-ns level time resolution enables separation of hits from different collisions when HHaS runs at an ultra-high event rate of around 100 MHz. The measurement of energy deposition helps distinguishing different nuclei with the same charge over mass ratio, which is crucial for hypernucleus and nuclear matter studies.
@@
HHaS accepts particles with transverse momenta above 50 MeV/c from 10 to 100 degrees in polar angle with full azimuthal coverage. It can clearly identify and measure $\pi^{\pm}$, $K^{\pm}$, $p$, $\bar{p}$, d, ${}^{3}He$, ${}^{4}He$, ${}^{6}Li$, $e^{\pm}$ and $\gamma$. Most impressively, the system can sustain interaction rates of $\sim$100 MHz for proton beam experiments and $\sim$1 MHz for heavy ion experiments, a capability that sets it apart from many existing experimental set-ups. These make HHaS a versatile spectrometer capable of many physics studies to be discussed in the following subsections.
@@ II.2 Search for Beyond-Standard-Model Physics and Test of Fundamental Symmetries with $\eta$ Meson Decays
The search for Beyond-Standard-Model (BSM) physics is one of the most exciting frontiers of particle physics. While the Standard Model (SM) has been extremely successful in explaining the fundamental particles and interactions observed so far, there remain unanswered questions, particularly concerning dark matter, neutrino masses, and violations of fundamental symmetries.
The $\eta$ meson, a neutral pseudoscalar meson, is an excellent candidate for probing BSM physics. Its decay channels are sensitive to new particles and interactions that are not predicted by the SM, offer potential avenues to search for dark portal particles like dark photons, dark scalars, and axion-like particles. Additionally, $\eta$ meson decays can help test discrete symmetry violations, such as CP and lepton flavor violations, providing crucial insights into fundamental physics.
@@
The combination of advanced spectrometer, high luminosity beams for large $\eta$ meson event collection, low background conditions, and the ability to accurately measure rare decays makes HHaS@HIAF an ideal facility for probing dark matter candidates. This is justified by simulation studies which shows that with HHaS at HIAF, more than one order of magnitude better sensitivities can be achieved for dark photon and dark scaler searches within certain mass ranges, and the CP violation test is two orders of magnitude more precise than similar existing results [5]. With the completion of the HHaS, these studies are set to open new avenues of research, contributing to the global effort to uncover physics beyond the Standard Model.
@@ II.3 Light Hadron Physics
Understanding how quarks combine into hadrons remains a central challenge in modern physics. While quantum chromodynamics (QCD) governs the strong interaction, its nonperturbative nature at low energies makes it difficult to tackle, leaving the strong force the least understood among the four fundamental interactions. The hadron spectrum serves as a crucial probe, driving major experimental efforts at BESIII, Belle II, and LHCb. High-precision measurements are essential for progress, and advanced facilities like HIAF and the upcoming CNUF are expected to make significant contributions. HHaS@HIAF’s broad energy coverage makes it an ideal platform for light hadron studies, allowing comprehensive investigations of hadronic interactions.
@@
These facilities provide high-density proton beams, potentially including polarized beams, which enable efficient production and investigation of light baryon resonances. This capability offers a robust platform for investigating the ’missing resonance’ problem. While nucleon resonances have been widely studied, multi-strange hyperons remain a greater challenge. Despite their discovery in the 1950s-60s, the Particle Data Group (PDG) confirms only a fraction of predicted states: of 23 $\Lambda$ states, only 14 are established; $\Sigma$ and $\Xi$ have 9 and 11, respectively; and $\Omega$ has just three. Moreover, key properties such as spin-parity assignments remain undetermined for many $\Xi$ and $\Omega$ states, even though lattice QCD and quark models predict around 70 states in this region [6]. In HAIF, hyperons can be produced via the kaon production process using the proton beams. The resulting secondary kaon beam offers a more efficient mechanism for hyperon production. With multi-kaon production, the scarcely produced $\Xi^{*}$ and $\Omega^{*}$ resonances—typically characterized by very low yields—can be effectively generated. Hence, the combination of high-density proton beams and secondary kaon beams provides significant advantages for investigating these elusive states.
@@
Thanks to its high precision, the HHaS@HIAF offers an excellent opportunity to study light exotic hadrons and probe anomalies such as the mass flip of N(1535) and Roper resonance. This anomaly is believed to arise from significant five-quark admixtures in its structure, challenging the conventional three-quark picture [7]. Furthermore, the internal structure of $\Lambda(1405)$ remains a hotly debated topic [6]. Competing models suggest it might be either a conventional three-quark state or a molecular-like state formed by strong $\bar{K}N$ interactions. Precise experimental data—especially on its line shape and decay patterns—are essential to resolve this issue, and HHaS@HIAF is well equipped to meet this challenge. In addition, the search for exotic states such as the dibaryon $d^{*}(2380)$ continues to be a major objective. Although initial hints have emerged from experiments like WASA@COSY, independent confirmation is still lacking [8]. With its high-precision proton beams, HHaS@HIAF holds great potential for delivering definitive insights into these elusive states and for uncovering further dibaryon resonances, deepening our understanding of the strong interaction in the nonperturbative regime.
@@ II.4 Nuclear Matter at High Baryon Density
Consequently, the energy ranges covered by HIAF and CNUF ($E_{k}<$ 9.1 GeV/u, $\sqrt{s_{NN}}<$ 4.5 GeV) represent a crucial region for searching the phase transition critical point. With its MHz-level event rate capability for heavy-ion collisions, large acceptance and outstanding particle identification, HHaS offers unique advantages for measuring critical fluctuation observables - including higher-order moments of conserved charges and light nucleus yield ratios - that are theoretically predicted to be sensitive to the critical point.
@@ II.5 Hypernucleus Physics
Heavy-ion collisions at several GeV energy can create a lot of hyperons, which may coalesce with nucleons in the collision system and form hypernuclei. Using heavy-ion beams with energies up to 9.1 GeV/u at HIAF, light hypernuclei can be created and studied. The good hit resolution of HHaS provides exceptional vertex resolution, which is crucial for eliminating almost all combinatorial backgrounds when reconstructing hypernuclei. HHaS’ 5-dimensional pixel tracking detector can also measure the energy loss of charged particles, which can be used to distinguish different nuclei with the same charge over mass ratio, such as d, 4He and 6Li. This is important for the identification of the daughter particles of hypernuclei. Taking advantage of the high luminosity of HIAF and the unprecedented high event rate of HHaS, various properties of hypernuclei, such as lifetime, mass, decay branching fractions, as well as their production yield, transverse momentum spectra and azimuthal anisotropy can be measured precisely, and new (multi-strange) hypernuclei can also be searched.
@@ III Muon Physics
Muon ($\mu^{-}$), as well as its antiparticle ($\mu^{+}$), has spin $1/2$ and a rest mass of 105.658 MeV/$c^{2}$. The lifetime of muon is about $2.2~{}\mu s$, making it the lightest unstable charged lepton. Taking this advantage, muon is regard as a promising probe for the new physics beyond the Standard Model (BSM). Moreover, it has broad applications in various areas of material science [13]. However, there is no operating accelerator muon source in China at present. The proposed muon beamlines based on the large scientific facilities under construction, HIAF and CiADS, are expected to break through the bottleneck [14, 15]. Some representative topics can be brought up in the context of the forthcoming muon beamlines in China.
@@ III.1 Charged lepton flavor violation
In the Standard Model, lepton flavor is a strictly conserved quantum number. The discovery of neutrino oscillations, which confirms that neutrinos have mass, provides the first direct evidence for physics beyond the Standard Model (BSM), demonstrating that lepton flavor can be violated in the neutral sector. Consequently, there is strong motivation to pursue experimental searches for BSM physics via charged lepton flavor violation (cLFV) processes. Popular channels include $\mu^{+}\rightarrow e^{+}\gamma$, $\mu^{+}\rightarrow e^{+}e^{-}e^{+}$, and $\mu^{-}N\rightarrow e^{-}N$ [16]. In recent years, notable experiments searching for these cLFV channels include MEG and its upgrade, MEG-II [17], Mu3e [18] at Paul Scherrer Institute, COMET in Japan [19, 20], and Mu2e in the USA [21]. Due to the lack of a dedicated muon beamline, no muon cLFV experiments have been proposed in China thus far. However, with the ongoing construction of the Huizhou large scientific facility, the research and development of muon sources is now underway. It paves the way for exploring unique muon cLFV processes in the coming years.
@@
In the baseline run plan, MACE will utilize a muon beamline with $10^{8}$ s-1 muon-on-target ($\mu$OT) flux, beam momentum of 24 MeV/$c$ and an root mean square (RMS) spread of 1.35 MeV/$c$. The data acquisition duration will be set at 1 year. As a result, MACE is expected to collect $1.20\times 10^{14}$ muonium decay events in the vacuum of interest. With the current MACE detector design and run plan, if no signal events are observed, it is expected to achieve a conversion probability upper bound, at 90% confidence level, of
$P({\text{M}}\to{\overline{\text{M}}})\lesssim 3.8\times 10^{-13}~{},$ (1)
which suggests an more than 2 orders of magnitude improvement in sensitivity compared to the MACS experiment at PSI. SMEFT calculations demonstrates that MACE has the potential to probe new physics at energy scales comparable to or beyond those of $\mu$TRISTAN, a future muon collider [36, 34].
@@
The Crystal Box experiment at the Los Alamos Meson Physics Facility (LAMPF) was the latest to constrain the branching ratio of $\mathcal{BR}(\mu^{+}\to e^{+}\gamma\gamma)<7.2\times 10^{-11}$ at the 90% confidence level in 1988 [42]. The latest experiment searching for the muonium annihilation $\mathrm{M}\to\gamma\gamma$ was conducted by C. M. York et al. in 1959 [43]. The experiment merely setup a loose bound of $\mathcal{O}(10^{-5})$ on $\mu^{+}e^{-}\to\gamma\gamma$ [43, 44]. No updated results or dedicated experimental plans have been proposed for these channels, as the limited geometrical acceptance of the current MEG calorimeter constrains its capability to probe the aforementioned processes. With the optimized detector system in MACE Phase-I, an improvement in sensitivity by one or more orders of magnitude is expected for the two decay modes mentioned above.
@@
where $f_{\text{cap}}$ is the muon capture efficiency at the muon stopping target, $f_{\text{geo}}$ is the geometrical acceptance of the detector. Take the proton energy is 1000 MeV at future CiADS, the corresponding stopped muon per proton is about ${}1\times 10^{-4}$. Assuming similar muon stopping target material and detector geometry as the Fermilab Mu2e-II proposal[45], it can be easily estimated that single event sensitivity can reach the level of $\sim 7\times 10^{-19}$.
@@ III.2 High-Precision Spectroscopy of Muonic Atoms
Testing QED in extreme fields using muonic atoms.
@@
It is well known that the effect of vacuum polarization is particularly significant in quantum electrodynamic effects on muonic atoms. With the development of superconducting transition edge sensor (TES) microcalorimeter, a proof-of-principle experiment for testing strong-field QED has been performed by measuring the X-ray spectroscopy of muonic neon atoms at J-PARC [55]. A high-intensity muon source driven by a continuous-wave superconducting linac is planned at CiADS which will significantly advance muonic atom x-ray spectroscopy [15]. We plan to employ the latest developed TES microcalorimeter in the energy range from 10 keV to 200 keV, [56], which is developed at the institute of Modern physics, to conduct precision X-ray measurements of muonic atoms for testing the QED effect under extreme strong electromagnetic field conditions. In the X-ray muonic atom experiment, a low-density gas target will be required to avoid rapid refilling of electrons into this system from surrounding atoms, as described in [54, 55].
Study of nuclear properties with muonic atoms.
@@
Significant progress has been achieved through several pioneering experimental programs. The CREMA collaboration extracted precise charge radii of 1,2H and 3,4He through laser spectroscopy of $\mu^{1,2}$, and $\mu^{3,4}$e+ [57, 63, 64, 65, 66] and is now extending these studies to muonic lithium. The $\mu$-X experiment focuses on high-resolution muonic atom X-ray spectroscopy of radioactive or rare isotopes using microgram target materials, both extracting nuclear charge radii and searching for charge-parity violation signals [67, 68, 69]. Complementarily, the QUARTET collaboration employs metallic magnetic calorimeters for high-precision X-ray spectroscopy of light muonic atoms, targeting charge radii determinations from Li to Ne isotopes with unprecedented accuracy [70, 71].
@@
The recently planned high-intensity muon beam at CiADS [15] will provide a significantly advance by measuring X-rays in higher-Z muonic atoms to investigate nuclear properties. The energy resolution of X-ray spectroscopy for muonic atoms will be improved with a factor of 5-10 by employing the advanced cryogenic TES and metallic magnetic cryogenic (MMC) micro-calorimeter technology as compared to conventional semiconductor detectors. Therefore, the nuclear properties, such as nuclear charge radii, magnetic dipole moments, and Zemach radii, can be investigated with a high precision across a broader range of isotopes. These measurements will provide unique experimental conditions for studying the properties of atomic nuclei and critical benchmarks for testing ab initio nuclear structure calculations. Taking full advantages of the current advanced experimental techniques and potential muonic atom capabilities at new facilities, the precision of charge radii measurements for stable nuclei across the nuclear chart could be significantly improved. This advancement would undoubtedly benefit charge radii studies of unstable nuclei as well.
Facilities at J-PARC, RIKEN-RAL, and PSI have concentrated on measuring the hyperfine splitting in $\mu$H to extract the proton Zemach radius—a parameter characterizing the convolution of proton’s charge and magnetization distributions that provides critical insights into nuclear spin structure [75, 76, 77]. The planned high-intensity muon beam at CiADS would enable hyperfine structure determinations with relative precision of 0.1–1 ppm via munoic atom X-ray. Priority targets will include muonic Li and Be isotopes, which allow high-precision benchmarking with the Zemach radii determined through spectroscopy of their electronic counterparts [78, 79, 80]. These measurements are essential for determining nuclear Zemach radii across a broader range of isotopes. CiADS also will significantly advance muonic atom spectroscopy by extending precision measurements to hyperfine splitting in higher-$Z$ muonic atoms. Implementing the advanced MMC micro-calorimeter, these measurements will provide critical benchmarks for testing ab initio nuclear structure calculations, particularly for nuclear polarizability contributions to muonic atom spectra that currently limit the precision in the determination of nuclear radii from muonic atom spectroscopy.
III.3 Probing and Knocking with Muons
Muons serve as valuable probes bridging fundamental physics and applications. Both cosmic and accelerator-based muons enable exploration beyond the Standard Model(SM), which, despite its success, cannot account for neutrino masses or dark matter(DM). The failure to detect Weakly Interacting Massive Particles(WIMPs) has spurred interest in alternative models, including low-mass or muon-philic DM [81, 82, 83].

### v2

@@ I Introduction
Table 1: Typical beam parameters for the HIAF-BRing
Ions Energy (GeV/u) Beam Intensity (ppp)
$^{238}\rm{U}^{35+}$ 0.835 $2.0\times 10^{11}$
$^{209}\rm{Bi}^{31+}$ 0.85 $2.4\times 10^{11}$
$^{78}\rm{Kr}^{19+}$ 1.7 $6.0\times 10^{11}$
$^{12}\rm{C}^{6+}$ 2.6 $1.2\times 10^{12}$
proton 9.3 $6.0\times 10^{12}$
Table 2: Proton beam parameters of CiADS Linac
Parameters Design 2028 2030 Upgrade ability
Energy (MeV) 500 600 600 1500
Current (mA) 5 0.5 5 10
Power (MW) 2.5 0.3 3 15
@@ II Huizhou Hadron Spectrometer
HIAF can provide full ion beam species ranging from protons to $^{238}\rm{U}$, with proton beam energy up to 9.3 GeV or $^{238}\rm{U}^{76+}$ beams up to 2.45 GeV/u. For its upgrade to CNUF, the facility can accelerate the proton beam to an energy of 25 GeV or $^{238}\rm{U}^{76+}$ beam to 7.3 GeV/u. This provides good opportunities for a wide range of physical research, including the search for new particles and new interactions beyond the standard model, the test of basic symmetries, the search for new (exotic) hadron states, exotic di-baryons and new (multi-strange) hypernuclei, precise measurement of hadron and hypernucleus prop- erties, the search and location of the nuclear matter phase boundary and critical point, etc. In order to conduct these physics studies, we plan to build an experimental spectrometer at the HIAF high energy terminal, called the Huizhou Hadron Spetrometer (HHaS). If acquiring sufficient funding support, we plan to take about 3 years to develop the key detector technologies for HHaS and another about 5 years to construct the whole spectrometer. Successful construction of HHaS will promote the development of medium- and high-energy particle physics and nuclear physics researches based on facilities in China.
@@ II.1 Conceptual Design
The 5-dimensional tracking system features an all-silicon pixel detector with $\sim$100 µm pitch pixels, incorporating both energy deposition and time measurement capabilities on each pixel. With $\sim<$20-µm hit position resolution and a 1.6-Tesla magnetic field, $\sim$3% momentum resolution can be achieved for particle p/m>p/m>1 with a tracker radius of merely 30 cm. This makes HHaS a very compact and cost-effective spectrometer. The good hit resolution also provides exceptional vertex resolution, which is crucial for clear reconstruction of strange particles and hypernuclei. The 10-ns level time resolution enables separation of hits from different collisions when HHaS runs at an ultra-high event rate of around 100 MHz. The measurement of energy deposition helps distinguishing different nuclei with the same charge over mass ratio, which is crucial for hypernucleus and nuclear matter studies.
@@
HHaS accepts particles with transverse momenta above 50 MeV/c from 10 to 100 degrees in polar angle with full azimuthal coverage. It can clearly identify and measure $\pi^{\pm}$, $K^{\pm}$, $p$, $\bar{p}$, d, 3He, 4He, 6Li, $e^{\pm}$ and $\gamma$. Most impressively, the system can record interaction events with rates of $\sim$100 MHz for proton beam experiments and $\sim$1 MHz for heavy ion experiments, a capability that sets it apart from many existing experimental set-ups. For example, the typical event rates for ATLAS, CMS, ALICE, STAR and HADES experiments are 75 kHz, 100 kHz, 50 kHz, 1 kHz and 20 kHz, respectively. These comprehensive abilities make HHaS a versatile spectrometer capable of many physics studies to be discussed in the following subsections.
@@ II.2 Search for Beyond-Standard-Model Physics and Test of Fundamental Symmetries with $\eta$ Meson Decays
The search for Beyond-Standard-Model (BSM) physics is one of the most exciting frontiers of particle physics. While the Standard Model (SM) has been extremely successful in explaining the fundamental particles and interactions observed so far, there remain unanswered questions, particularly concerning dark matter, neutrino masses, and violations of fundamental symmetries [2, 3, 4].
The $\eta$ meson, a neutral pseudoscalar meson, is an excellent candidate for probing BSM physics particularly through its deacy channels. The physics potential of $\eta$ mesons in BSM searches has been explored extensively, especially in the context of dark matter, neutrino masses, and discrete symmetry violations. Theoretical studies, including those from the Bonn/Julich group [5, 6], predict that $\eta$ meson decays can be sensitive to new particles and interactions not predicted by the SM, such as dark photons, axion like particles, and dark scalars [7]. The decay modes are particularly valuable for searching for dark matter candidates and testing CP violations and lepton flavor symmetries.
@@
The combination of advanced spectrometer, high luminosity beams for large $\eta$ meson event collection, low background conditions, and the ability to accurately measure rare decays makes HHaS@HIAF an ideal facility for probing dark matter candidates. This is justified by simulation studies which shows that with HHaS at HIAF, more than one order of magnitude higher sensitivities can be achieved for dark photon and dark scaler searches within certain mass ranges, and the CP violation test is two orders of magnitude more precise than similar existing results [11]. With the completion of the HHaS, these studies are set to open new avenues of research, contributing to the global effort to uncover physics beyond the SM.
@@ II.3 Light Hadron Physics
Understanding how quarks combine into hadrons remains a central challenge in modern physics. While quantum chromodynamics (QCD) governs the strong interaction, its nonperturbative nature at low energies makes it difficult to tackle, leaving the strong force the least understood among the four fundamental interactions. The hadron spectrum serves as a crucial probe, driving major experimental efforts at BES III, Belle II, LHCb and J-PARC experiments. High-precision measurements are essential for progress, and advanced facilities like HIAF and the upcoming CNUF are expected to make significant contributions. HHaS@HIAF’s broad energy coverage makes it an ideal platform for light hadron studies, allowing comprehensive investigations of hadronic interactions.
@@
These facilities provide high-intensity proton beams, potentially including polarized beams, which enable efficient production and investigation of light baryon resonances. This capability offers a robust platform for investigating the “missing resonance” problem. While nucleon resonances have been widely studied, multi-strange hyperons remain a greater challenge. Despite their discovery in the 1950s-60s, the Particle Data Group (PDG) confirms only a fraction of predicted states: of 23 $\Lambda$ states, only 14 are established; $\Sigma$ and $\Xi$ have 9 and 11, respectively; and $\Omega$ has just three. Moreover, key properties such as spin-parity assignments remain undetermined for many $\Xi$ and $\Omega$ states, even though lattice QCD and quark models predict around 70 states in this region [12]. Usually, the yield of $\Xi^{*}$ and $\Omega^{*}$ resonance states is extremely low. The high-intensity potentially-polarized proton beam at HIAF and the unprecedented high event rate of HHaS enable effective production and detection of these states. In addition, the potential secondary kaon beam at HIAF also provides an efficient mechanism to produce hyperons, which is similar to the J-PARC secondary kaon beam experiments. Therefore, the high-intensity proton beam on HIAF and the secondary kaon beam that may be produced in the future bring significant advantages to the study of these difficult-to-detect states.
@@
HHaS@HIAF also offers an excellent opportunity to study light exotic hadrons and probe anomalies such as the mass flip of NN(1535) and Roper resonance. This anomaly is believed to arise from significant five-quark admixtures in its structure, challenging the conventional three-quark picture [13]. Furthermore, the internal structure of $\Lambda(1405)$ remains a hotly debated topic [12]. Competing models suggest it might be either a conventional three-quark state or a molecular-like state formed by strong $\bar{K}N$ interactions. Precise experimental data—especially on its line shape and decay patterns—are essential to resolve this issue, and HHaS@HIAF is well equipped to meet this challenge. In addition, the search for exotic states such as the dibaryon $d^{*}(2380)$ continues to be a major objective. Although initial hints have emerged from experiments like WASA@COSY, independent confirmation is still lacking [14]. With its high-intensity proton beams, HHaS@HIAF holds great potential for delivering definitive insights into these elusive states and for uncovering further dibaryon resonances, deepening our understanding of the strong interaction in the nonperturbative regime.
@@ II.4 Nuclear Matter at High Baryon Density
Consequently, the energy ranges covered by HIAF and CNUF ($E_{k}<$ 9.1 GeV/u, $\sqrt{s_{NN}}<$ 4.5 GeV) represent a crucial region for searching the phase transition critical point. With its MHz-level event rate capability for heavy-ion collisions, large acceptance and good particle identification ability, HHaS offers unique advantages for measuring critical fluctuation observables - including higher-order moments of conserved charges and light nucleus yield ratios - that are predicted to be sensitive to the critical point.
@@ II.5 Hypernucleus Physics
Heavy-ion collisions at several GeV energy can create a lot of hyperons, which may coalesce with nucleons in the collision system and form hypernuclei. STAR@RHIC, ALICE@LHC and future CBM@FAIR uses or plans to use heavy-ion collisions to study light hypernuclei. Using heavy-ion beams with energies up to 9.1 GeV/u at HIAF, light hypernuclei can be created abundantly and studied. The good hit resolution of HHaS provides exceptional vertex resolution, which is crucial for eliminating almost all combinatorial backgrounds when reconstructing hypernuclei. HHaS’ 5-dimensional pixel tracking detector can also measure the energy loss of charged particles, which can be used to distinguish different nuclei with the same charge over mass ratio, such as d, 4He and 6Li. This is important for the identification of the daughter particles of hypernuclei. Taking advantage of the high luminosity of HIAF and the unprecedented event rate of HHaS, various properties of hypernuclei, such as lifetime, mass, decay branching fractions, as well as their production yield, transverse momentum spectra and azimuthal anisotropy can be measured precisely, and new (multi-strange) hypernuclei can also be searched.
@@ III Muon Physics
[…]
@@ III.1 Charged lepton flavor violation
In the SM, lepton flavor is a strictly conserved quantum number. The discovery of neutrino oscillations, which confirms that neutrinos have mass, provides the first direct evidence for physics BSM, demonstrating that neutral leptons could exhibit flavor mixing effects. Consequently, there is strong motivation to pursue experimental searches for BSM physics via charged lepton flavor violation (cLFV) processes. Popular channels include $\mu^{+}\rightarrow e^{+}\gamma$, $\mu^{+}\rightarrow e^{+}e^{-}e^{+}$, and $\mu^{-}N\rightarrow e^{-}N$ [22]. In recent years, notable experiments searching for these cLFV channels include MEG and its upgrade, MEG-II [23], Mu3e [24] at Paul Scherrer Institute, COMET in Japan [25, 26], and Mu2e in the USA [27]. Due to the lack of a dedicated muon beamline, no muon cLFV experiments have been proposed in China thus far. However, with the ongoing construction of the Huizhou large scientific facility, the research and development of muon sources is now underway. It paves the way for exploring unique muon cLFV processes in the coming years.
@@
[…]
@@
The Crystal Box experiment at the Los Alamos Meson Physics Facility (LAMPF) was the latest to constrain the branching ratio of $\mathcal{BR}(\mu^{+}\to e^{+}\gamma\gamma)<7.2\times 10^{-11}$ at the 90% confidence level in 1988 [48]. The latest experiment searching for the muonium annihilation $\mathrm{M}\to\gamma\gamma$ was conducted by C. M. York et al. in 1959 [49]. The experiment merely setup a loose bound of $\mathcal{O}(10^{-5})$ on $\mu^{+}e^{-}\to\gamma\gamma$ [49, 50]. No updated results or dedicated experimental plans have been proposed for these channels, as the limited geometrical acceptance of the current MEG calorimeter constrains its capability to probe the aforementioned processes. With the optimized detector system in MACE Phase-I, sensitivities of $\mathcal{O}(10^{-12}-10^{-13})$ are expected for the two decay modes mentioned above.
@@
where $f_{\text{cap}}$ is the muon capture efficiency at the muon stopping target, $f_{\text{geo}}$ is the geometrical acceptance of the detector. For the muon LFV program, current experiments face limitations primarily in two areas: statistics and systematics, respectively: 1. Experiments such as MEGII and Mu3e are primarily limited by accidental coincidence backgrounds, which are a form of systematic error. Consequently, simply increasing the beam intensity (to improve statistics) will not substantially enhance their sensitivity or improve the upper limits. 2. In contrast, experiments like COMET and Mu2e [51] are mainly constrained by the available muon flux, representing statistical errors. Since these setups do not suffer from significant accidental backgrounds, systematic uncertainties can be minimized through careful facility design and precise simulations. Therefore, increasing the muon beam intensity is expected to significantly improve their sensitivity. Looking ahead, more intense muon fluxes are anticipated from upcoming accelerator facilities in China. By applying the same physics processes–such as muon-to-electron conversion used in COMET and Mu2e–extrapolations based on projected beam intensities are both ambitious and plausible. These developments hold the potential to push the boundaries of LFV searches to new, more stringent limits.
@@ III.2 High-Precision Spectroscopy of Muonic Atoms
III.2.1 Testing QED in extreme fields using muonic atoms.
@@
It is well known that the effect of vacuum polarization is particularly significant in quantum electrodynamic effects on muonic atoms. With the development of superconducting transition edge sensor (TES) microcalorimeter, a proof-of-principle experiment for testing strong-field QED has been performed by measuring the X-ray spectroscopy of muonic neon atoms at J-PARC [61]. A high-intensity muon source driven by a continuous-wave superconducting linac is planned at CiADS which will significantly advance muonic atom x-ray spectroscopy [21]. We plan to employ the latest developed TES microcalorimeter in the energy range from 10 keV to 200 keV [62], which is developed at the institute of Modern physics, to conduct precision X-ray measurements of muonic atoms for testing the QED effect under extreme strong electromagnetic field conditions. In the X-ray muonic atom experiment, a low-density gas target will be required to avoid rapid refilling of electrons into this system from surrounding atoms, as described in [60, 61].
III.2.2 Study of nuclear properties with muonic atoms.
@@
Significant progress has been achieved through several pioneering experimental programs. The CREMA collaboration extracted precise charge radii of 1,2H and 3,4He through laser spectroscopy of $\mu^{1,2}$H, and $\mu^{3,4}$He+ [63, 69, 70, 71, 72] and is now extending these studies to muonic lithium. The $\mu$-X experiment focuses on high-resolution muonic atom X-ray spectroscopy of radioactive or rare isotopes using microgram target materials, both extracting nuclear charge radii and searching for charge-parity violation signals [73, 74, 75]. Complementarily, the QUARTET collaboration employs metallic magnetic calorimeters (MMC) for high-precision X-ray spectroscopy of light muonic atoms, targeting charge radii determinations from Li to Ne isotopes with unprecedented accuracy [76, 77].
@@
The recently planned high-intensity muon beam at CiADS [21] will provide a significantly advance by measuring X-rays in higher-Z muonic atoms to investigate nuclear properties. The energy resolution of X-ray spectroscopy for muonic atoms will be improved with a factor of 5-10 by employing the advanced cryogenic TES and MMC micro-calorimeter technology as compared to conventional semiconductor detectors. Therefore, the nuclear properties, such as nuclear charge radii, magnetic dipole moments, and Zemach radii, can be investigated with a high precision across a broader range of isotopes. These measurements will provide unique experimental conditions for studying the properties of atomic nuclei and critical benchmarks for testing ab initio nuclear structure calculations. Taking full advantages of the current advanced experimental techniques and potential muonic atom capabilities at new facilities, the precision of charge radii measurements for stable nuclei across the nuclear chart could be significantly improved. This advancement would undoubtedly benefit charge radii studies of unstable nuclei as well.
[…]
III.3 Probing and Knocking with Muons(PKMu)
Muons serve as valuable probes bridging fundamental physics and applications. Both cosmic and accelerator-based muons enable exploration beyond the Standard Model(SM), which, despite its success, cannot account for neutrino masses or dark matter(DM). The absence of any direct detection of Weakly Interacting Massive Particles (WIMPs) to date has prompted the scientific community to explore alternative models, including light-mass or muon-philic DM. [87, 88, 89].